# Enhancing Wrist Abnormality Detection with YOLO: Analysis of State-of-the-art Single-stage Detection Models

### v1

@@ 3.4.3 The YOLOv7 Model
YOLOv7 also has three main components discussed before with several changes. The E-ELAN is a component in YOLOv7 that uses expand, shuffle, and merge cardinality to continuously improve network learning without disrupting the gradient path. Wang et al. (2022). Other notable changes include Model Scaling techniques, Re-parameterization planning, and Auxiliary Head Coarse-to-Fine. Model scaling is a technique used to adapt key characteristics of a model to align with specific application requirements. This includes adjusting the width (number of channels), depth (number of stages), and resolution (input image size) of the model. The scaling of object detection models requires knowledge of the network depth, width, and resolution on which it is trained.
YOLOv7 utilizes a compound scaling technique, which simultaneously scales the depth and width of the network by concatenating layers. This method has been shown through ablation studies to maintain optimal model architecture while scaling for different sizes. Without this technique, an increase in depth alone may cause a decrease in hardware efficiency due to a change in the ratio between input and output channels of a transition layer. YOLOv7’s compound scaling technique prevents such negative effects on performance. Re-parameterization techniques aim to create a more robust model by averaging a set of weights. Recent research has focused on module-level re-parameterization, where specific parts of the network are targeted. YOLOv7 uses gradient flow propagation to determine which modules should be re-parameterized. The YOLO network head generates the final predictions, however, an auxiliary head located in the middle of the network can be beneficial during training. The auxiliary head is supervised along with the final head. However, it does not train as efficiently because it is closer to the prediction, thus the YOLOv7 authors experimented with different levels of supervision for the auxiliary head, using a coarse-to-fine approach where supervision is passed back from the final head at various granularities.
The YOLOv7 model comprises of seven different variants, which include "P5" models (v7, v7x, and v7-tiny) and "P6" models (d6, e6, w6, and e6e). These variants are compound-scaled versions of the same architecture, each of which offers a different level of detection accuracy and performance when trained on the standard COCO dataset. This variation in performance is illustrated in Fig. 5.
In the experimentation of these 7 variants of the YOLOv7 model, all variants underwent a training phase with a duration of 100 epochs from scratch with 16 samples as batch size. The standard hyperparameters were applied and the optimization algorithm employed was Stochastic Gradient Descent (SGD) with an initial learning rate $\alpha=1\times 10^{-2}$, final learning rate $\alpha_{f}=1\times 10^{-1}$, momentum = 0.937, weight decay = $5\times 10^{-4}$, warmup momentum = 0.8, and warmup bias lr = 0.1. The "P5" models within YOLOv7, namely, v7, v7x, and v7-tiny were trained with an input resolution of 640 pixels, while the "P6" models, namely, d6, e6, w6, and e6e were trained with an input resolution of 448 pixels due to computational constraints, and although it may have a negative effect on the model’s performance, this issue is compensated for by utilizing mosaic augmentation within YOLOv7. Mosaic augmentation is a technique used to increase the diversity of training data by combining multiple small images into a larger "mosaic" image. This can help to improve the robustness of object detection models by exposing them to a wider variety of object scales, orientations, and backgrounds. During the evaluation phase, all variants were tested on a test set of 1016 samples, using an Intersection over Union (IoU) threshold of 0.5.
Figure 5: Performance comparison of YOLOv7 variants on COCO Wang et al. (2022).
@@ 3.4.4 The YOLOv8 Model
YOLOv8 is reported to provide significant advancements in object detection as well as image segmentation when compared to previous YOLO models, particularly in compact versions that are implemented on less powerful hardware. At the time of writing this paper, the architecture of YOLOv8 is not fully disclosed and some of its features are still under development. As of now, it’s been confirmed that the system has a new backbone, uses an anchor-free design, has a revamped detection head, and has a newly implemented loss function. We have included the performance of this model on the GRAZPEDWRI-DX dataset as a benchmark for future studies, as further improvements to YOLOv8 may surpass the results obtained in this study.
YOLOv8 comes in five versions at the time of release (January 10, 2023), namely, "n", "s", "m", "l", and "x". The performance of these variants on a COCO dataset is shown in Fig. 6. Just as before, all variants underwent 100 epochs of training from scratch, with standard hyperparameters. The image resolution was set at 640 pixels, with a batch size of 16 samples. The optimization algorithm employed was Stochastic Gradient Descent (SGD) as before with a starting learning rate of $1\times 10^{-2}$, final learning rate of $1\times 10^{-2}$, the momentum of 0.937, weight decay of $5\times 10^{-4}$, warmup momentum of 0.8, and warmup bias lr of 0.1. As before, all variants were tested on a test set of 1016 samples, using an IoU threshold of 0.5.
Figure 6: Performance comparison of YOLOv8 variants on COCO ultralytics (2023).
@@ 3.4.5 Faster R-CNN
The Faster R-CNN model consists of three main components: the backbone, the region proposal network (RPN), and a detection network. In this study, a ResNet50 backbone with FPN was used for feature extraction from the input image. Anchors were generated for each feature, and a set of anchor boxes with variable sizes and aspect ratios were created for each anchor. The RPN was responsible for selecting appropriate anchor boxes and passing them onto the next layer. The classifier within RPN predicted if an anchor box contained an object, determined by an IoU threshold of 0.5. The regressor within RPN predicted offsets for the anchor boxes containing objects to fit them tightly to the ground truth labels. Lastly, the RoI pooling layer converted variable-sized proposals to a fixed size to run a classifier and regress a bounding box. Fig. 7 illustrates the architecture of Faster R-CNN.
The Faster R-CNN model underwent 100 epochs with an image size of 640 pixels and 16 samples as batch size. The default parameters were used and the optimization algorithm employed was Stochastic Gradient Descent (SGD) with a learning rate of $\alpha=1\times 10^{-3}$, a momentum of 0.9, and weight decay of $5\times 10^{-4}$. It is important to note that, as with YOLO models, the selection of these parameters is not deliberate, they are the default settings. During the evaluation phase, each variant was tested on 1016 randomly selected samples, using an Intersection over Union (IoU) threshold of 0.5 for inference.
@@ 3.5 Evaluation Metrics: mAP
3.6 Supplementary Materials
The supplementary materials, including source code, and dataset split can be accessed through the following links:
@@ 5 Results & Discussion
Figures 12 and 13 provide an overview of the mAP scores obtained for fracture class as well as across all classes by all YOLO variants and Faster R-CNN. In applications where false positives are costly, a model with high precision may be preferable, while in situations where missing detections are costly, a model with high recall may be more desirable. The mean Average Precision (mAP) serves as a comprehensive measure of the model’s performance. Therefore, we selected the best-performing variant within each YOLO model based on the highest mAP achieved for the fracture class and overall performance across all classes. We have also compared their mAP scores to each other as well as with that of the Faster R-CNN model, as illustrated in Table 16. We also evaluated the performance of all variants, including Faster R-CNN, on a challenging image containing multiple objects of interest, including 2 fractures, 3 periosteal reactions, 1 metal, and 1 text. The bounding box estimates for these objects from each variant and Faster R-CNN are illustrated in Fig. 14.
@@
It is clear from Table 16 that the variant "YOLOv8x" of YOLOv8 is the best-performing variant out of all the variants employed in this study. The results presented in this study using the variant "YOLOv8x" represent a significant improvement upon the ones originally presented in Nagy et al. (2022) for the fracture class. In that paper, the model variant "YOLOv5m" trained on COCO weights achieved a mean average precision (mAP) score of 0.93 for fracture detection and an overall mAP score of 0.62 at an IoU threshold of 0.5. In contrast, the results obtained in this study demonstrate a higher mAP score of 0.95 for fracture detection and an overall mAP of 0.77 at an IoU threshold of 0.5. Fig. 10(a), 10(b), 10(c), and 10(d) present the F1 versus Confidence, Recall versus Confidence, Precision versus Confidence, and Precision versus Recall curves, respectively, for the variant "YOLOv8x" across all classes. These curves provide a visual representation of the model’s performance on different confidence intervals and allow for a more thorough evaluation of its capabilities. The F1 versus Confidence curve shows the relationship between the model’s F1 score, which is a measure of the balance between precision and recall, and the confidence of its predictions. The Recall versus Confidence curve illustrates the model’s ability to correctly identify objects, while the Precision versus Confidence curve demonstrates the proportion of correct predictions made by the model. The Precision versus Recall curve shows the trade-off between the model’s precision and recall, with higher precision typically corresponding to lower recall and vice versa. Additionally, a confusion matrix 11 is shown for the variant "YOLOv8x".

### v2

@@ 3.3.3 The YOLOv7 Model
YOLOv7 comes with several changes, including E-ELAN, which uses expand, shuffle, and merge cardinality to improve network learning without disrupting the gradient path. Other changes include Model Scaling techniques, Re-parameterization planning, and Auxiliary Head Coarse-to-Fine. Model scaling adjusts the width, depth, and resolution of a model to align with specific application requirements. YOLOv7 uses compound scaling to simultaneously scale network depth and width by concatenating layers, maintaining optimal architecture while scaling.
Re-parameterization techniques use gradient flow propagation to identify modules that require averaging weights for robustness. An auxiliary head in the middle of the network improves training but requires a coarse-to-fine approach for efficient supervision. The YOLOv7 model consists of seven variants: "P5" models (v7, v7x, and v7-tiny) and "P6" models (d6, e6, w6, and e6e).
@@ 3.3.4 The YOLOv8 Model
YOLOv8 is reported to provide significant advancements in object detection when compared to previous YOLO models, particularly in compact versions that are implemented on less powerful hardware. At the time of writing this paper, the architecture of YOLOv8 is not fully disclosed and some of its features are still under development. As of now, it’s been confirmed that the system has a new backbone, uses an anchor-free design, has a revamped detection head, and has a newly implemented loss function. We have included the performance of this model on the GRAZPEDWRI-DX dataset as a benchmark for future studies, as further improvements to YOLOv8 may surpass the results obtained in this study. YOLOv8 comes in five versions at the time of release (January 10, 2023), namely, "n", "s", "m", "l", and "x".
@@ 3.3.5 Faster R-CNN
The Faster R-CNN model includes a backbone, an RPN (regional proposal network), and a detection network. ResNet50 with FPN is used as the backbone for feature extraction. Anchors with variable sizes and aspect ratios are generated for each feature. The RPN selects appropriate anchor boxes using a classifier that predicts if an anchor box contains an object based on an IoU threshold of 0.5. The regressor predicts offsets for anchor boxes containing objects to fit them tightly to the ground truth labels. Finally, the RoI pooling layer converts variable-sized proposals to a fixed size to run a classifier and regress a bounding box. Fig. 3 illustrates the architecture of Faster R-CNN.
@@
3.4 Training Details
In the experimentation of YOLO variants, standard hyperparameters were utilized. The input resolution was fixed at 640 pixels. The optimization algorithm employed was SGD with an initial learning rate $\alpha=1\times 10^{-2}$, final learning rate $\alpha_{f}=1\times 10^{-2}$ (except for YOLOv7 variants with a final learning rate $\alpha_{f}=1\times 10^{-1}$), momentum = 0.937, weight decay = $5\times 10^{-4}$. Each variant/model underwent 100 epochs of training from scratch and was observed to converge between 90-100 epochs. Every variant was trained with a batch size of 16 except for the "P6" variants of YOLOv7 namely (d6, e6, w6, e6e) which were trained with a batch size of 8 due to computational constraints.
With Faster R-CNN, the only difference was the learning rate of $\alpha=1\times 10^{-3}$, momentum of 0.9 and weight decay of $5\times 10^{-4}$. All other parameters were the same as YOLO variants. As with YOLO models, the selection of these parameters is not deliberate, they are the default settings.
All binary classifiers were trained for a maximum of 100 epochs using a batch size of 64. The learning rate was set at $1\times 10^{-3}$. The Adam optimization algorithm guided the training process. Input images were standardized to a resolution of 224 pixels.
@@ 3.5 Evaluation Metrics: mAP
3.5.1 Sensitivity
Sensitivity, in the context of our model, pertains to its capacity to accurately recognize true detections among all positive detections within the dataset. Specifically, it gauges the model’s ability to correctly identify the presence of a fracture or abnormality. We prioritize this metric due to the potential consequences of false negatives in wrist trauma cases. Failure to detect fractures is a frequent reason for differences in diagnosis between the initial interpretation of X-ray images and the final analysis conducted by certified radiologists. The calculation for sensitivity is as follows:
Sensitivity=True PositivesTrue Positives+False $\text{Sensitivity}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (6)
@@ 5 Results & Discussion
Figures 8 and 9 provide an overview of the mAP scores obtained for fracture class as well as across all classes by all YOLO variants and Faster R-CNN. In applications where false positives are costly, a model with high precision may be preferable, while in situations where missing detections are costly, a model with high recall may be more desirable. The mean Average Precision (mAP) serves as a comprehensive measure of the model’s performance. Therefore, we selected the best-performing variant within each YOLO model based on the highest mAP achieved for the fracture class and overall performance across all classes. We have also compared their mAP scores to each other as well as with that of the Faster R-CNN model, as illustrated in Table 16. We also evaluated the performance of all variants, including Faster R-CNN, on a challenging image containing multiple objects of interest, including 2 fractures, 3 periosteal reactions, 1 metal, and 1 text. The bounding box estimates for these objects from each variant and Faster R-CNN are illustrated in Fig. 10.
@@
It is clear from Table 16 that the variant "YOLOv8x" of YOLOv8 is the best-performing variant out of all the variants employed in this study. The results presented in this study using the variant "YOLOv8x" represent a significant improvement upon the ones originally presented in Nagy et al. (2022) for the fracture class. In that paper, the model variant "YOLOv5m" trained on COCO weights achieved a mean average precision (mAP) score of 0.93 for fracture detection and an overall mAP score of 0.62 at an IoU threshold of 0.5. In contrast, the results obtained in this study demonstrate a higher mAP score of 0.95 for fracture detection and an overall mAP of 0.77 at an IoU threshold of 0.5. Fig. 6(a), 6(b), 6(c), and 6(d) present the F1 versus Confidence, Recall versus Confidence, Precision versus Confidence, and Precision versus Recall curves, respectively, for the variant "YOLOv8x" across all classes. These curves provide a visual representation of the model’s performance on different confidence intervals and allow for a more thorough evaluation of its capabilities. The F1 versus Confidence curve shows the relationship between the model’s F1 score, which is a measure of the balance between precision and recall, and the confidence of its predictions. The Recall versus Confidence curve illustrates the model’s ability to correctly identify objects, while the Precision versus Confidence curve demonstrates the proportion of correct predictions made by the model. The Precision versus Recall curve shows the trade-off between the model’s precision and recall, with higher precision typically corresponding to lower recall and vice versa. Additionally, a confusion matrix 7 is shown for the variant "YOLOv8x".